OpenStudy (anonymous):

Three of your friends, Anna, Blake, and Christine, run to you to settle a dispute. They were simplifying a radical expression into an exponential expression, but they reached different answers. Wisely, you decide to look at their work to see if you can spot the source of confusion. Explain to Anna, Blake, and Christine who is correct and identify any errors that you find. Provide the correct manner to fix those problems and identify the correct answer. Use complete sentences.

OpenStudy (anonymous):

What is their work?

OpenStudy (anonymous):

hold on I'm getting it

OpenStudy (anonymous):

Anna's work is correct. Blake subtracted 3 instead of multiplying 6 and x. Christine multiplied by 3 instead of 1/3.

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

Can you help me with one more problem @sidster182?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

What is the domain and range of f(x) = 75(1.72)?

OpenStudy (anonymous):

75(1.72) is a term. You can only find the domain and range of an equation.

OpenStudy (anonymous):

I thought that was an equation

OpenStudy (anonymous):

Usually equations contain a variable such as x. If that is the question given the domain would just be all real numbers and the range would be 129.

OpenStudy (anonymous):

how would the range be 129?

OpenStudy (anonymous):

75 multiplied by 1.72 gives you 129. The equation you're giving me to find the domain and range is basically y=129.

OpenStudy (anonymous):

but 75 is the y-intercept so shouldn't the range be anything greater than 75?

OpenStudy (anonymous):

is f(x)=75(1.72) the original equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

actually I did forget a little thing

OpenStudy (anonymous):

f(x) = 75(1.72)^x

OpenStudy (anonymous):

Okay, that makes a lot more sense. In this case your range would be all positive real numbers since the variable is in the exponent position.

OpenStudy (anonymous):

ok thank you that makes more sense to me too